Dec. 26 1999 Owl shift summary.

Valeri Jejer (jejer@uvahea.phys.virginia.edu)
Sun, 26 Dec 1999 08:10:54 -0500 (EST)

Shift Summary for owl shift, 26/Dec/99

On Shift: Ruth, Valeri.

Hours of beam available at all intensities: 6.85
Hours of beam available at nominal intensity: 6.85
Hours of E832 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

debugging trigger configuration with no fruits.

========================================================================
Runs ended this shift:
Run 15291 - - BEAM_799 tested configuration with no fruits in trigger
Run 15292 - in progress - BEAM_799 (no fruits)

Tapes brought to FCC this shift:

(including tapes for online split)

none
========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Smooth datataking at 7e12 with fruits out of readout and trigger after
some initial problems with DA and triggers.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15292 00:44 Start new run with no fruits (and fruit related YTF
and STT) in trigger. Had some problems starting run
including lost permits, busies from various crates,
multiple kfrends. Run started with triggers hanging
4-5 seconds before end of spill due to higher rates.

01:20 Sasha L. asked MCR to lower beam intensity to 7e12
and DAQ_DEAD errors disappeared at ~7.5e12.

40 01:40 Beginning of "stable" running conditions in the
new configuration. Turned off "SERIOUS ALARMS"
monitor's sound because L2/L1 scaler alarms are
going off every spill. Should be turned back on
after the limits are adjusted or fruits back in
readout whichever comes first.

03:08 DYC53 full followed by two kfrends.

03:22 Beam goes away.

126 03:32 Beam is back. SEED2 horizontal position is slightly
off for this and next spill (-4 - -5), as well as
for spills 129-130.

04:30 Noticed that CsI 983 has pedestal drifted by 7.5
counts.

04:55 Mounted UPF065 on ktev1.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Don't worry about the "heater" temp. Just be sure the "CsI-Bot" and
CsI-Top" readings are about 15 and 19 degrees, respectively.

-inefficient HCC bit 1767. Bad connection in ET that will be worked on
next long down time.
-latches are out of the readout so some of the monitoring histograms look
funny.

All fruits are out of trigger. Ignore Fastbus RR177 rack related SDAQ red
buttons.

Turn on sound on SERIOUS ALARMS monitor when L2/L1 scaler alarms
disappear.

CsI 983 has ped drift 7.5 counts.

CIA/SEM scaler alarm is going off every few spills.